OpenStudy (anonymous):

integrate (4x^2)/(x^3+2) where do I even start? I know part of it is going to be ln (x^3+2)...

OpenStudy (anonymous):

You get really lucky in this one, because you can do the \(u\) sub on the denominator.

OpenStudy (anonymous):

how would i know to do that?

OpenStudy (anonymous):

\[ d(x^3+2)=3x^2dx \]And so \[ \int \frac{4x^2}{x^3+2}dx = \int \frac{4}{3}\frac{1}{x^3+2}(3x^2dx)= \frac{4}{3}\int \frac{1}{x^3+2}d(x^3+2) \]

OpenStudy (anonymous):

I end up with 4/3 ln (x^3+2) but how do I know to u sub or try part or trig sub or something else?

OpenStudy (anonymous):

\[ \frac{4}{3}\int \frac{1}{x^3+2}d(x^3+2) \]Is equivalent to saying: \[ \frac{4}{3}\int \frac{1}{u}du \]where \(u = x^3+2\)

OpenStudy (anonymous):

There are many different things to rule out, so it is very complicated.

OpenStudy (anonymous):

Do \(u\) sub when you think you have a derivative of another function.

OpenStudy (anonymous):

=-).....hehe yup I think the complicated is what throws me off

OpenStudy (anonymous):

Look at the derivative of the denominator, or the derivative of the expression in the square root.

OpenStudy (anonymous):

If you see it, then you know to use \(u\) sub.

OpenStudy (anonymous):

If you have \[ \int \frac{c}{a\pm bx^2}dx \]Or \[ \int c\sqrt{a\pm bx^2}dx \]Or even \[ \int \frac{c}{\sqrt{a\pm bx^2}}dx \]You gotta use trig sub because the derivative of \(a\pm bx^2\) is \(\pm bx\) and you don't have an \(x\) to factor out.

OpenStudy (anonymous):

I guess what I really need is a little algorithm to apply when I'm first given the problems - if you pick the wrong integration method it is almost hopeless

OpenStudy (anonymous):

That algorithm would NOT be little.

OpenStudy (anonymous):

Just start out considering \(u\) sub. If you don't see any obvious derivatives, move on to the next method.

OpenStudy (anonymous):

hehe well guessing is not working for me...I tried to do it by parts when I first got it and a page and half later I gave up

OpenStudy (anonymous):

When you see those forms I just wrote a moment ago, you know immediately you need trig sub.

OpenStudy (anonymous):

Trig subs stick out like a sore thumb.

OpenStudy (anonymous):

kk I'll try to remember

OpenStudy (anonymous):

Parts... it not an easy method to master. In general you are looking for which part has an nice integral and which part has a nice derivative.

OpenStudy (anonymous):

When it comes to rational functions... that is polynomial over polynomial... Then: 1. Try to factor 2. Trig sub? 3. U sub If those don't work, you're kinda screwed.

OpenStudy (anonymous):

If something not working was an objective standard lol thanks for help I'll let you go help someone a little less incorrigible